# WisPaper: Your AI Scholar Search Engine

### v1

@@ Abstract
Researchers struggle to efficiently locate and manage relevant literature within the exponentially growing body of scientific publications. We present WisPaper, an intelligent academic retrieval and literature management platform that addresses this challenge through three integrated capabilities: (1) Scholar Search, featuring both quick keyword-based and deep agentic search modes for efficient paper discovery; (2) Library, a customizable knowledge base for systematic literature organization; and (3) AI Feeds, an intelligent recommendation system that automatically delivers relevant new publications based on user interests. Unlike existing academic tools, WisPaper provides a closed-loop workflow that seamlessly connects literature discovery, management, and continuous tracking of research frontiers. Our multilingual and multidisciplinary system significantly reduces the time researchers from diverse backgrounds spend on paper screening and management, enabling them to focus on their core research activities. The platform is publicly accessible and serves researchers across academia and industry.
@@ 1 Introduction
The exponential growth of scientific publications presents unprecedented challenges for researchers attempting to stay current in their fields [5]. Efficiently identifying relevant work and systematically managing findings over time have become increasingly time-consuming and cognitively demanding[12].
Current academic search systems exhibit significant limitations in supporting comprehensive research workflows [8]. Traditional platforms like Google Scholar [^1] https://scholar.google.com and Semantic Scholar [^2] https://www.semanticscholar.org/ rely primarily on keyword-based retrieval, which struggles with conceptual queries and semantic understanding. Recent advances have leveraged large language models (LLMs) [15, 3, 20] for query reformulation and expansion [10, 2, 14, 16], improving initial query understanding. Yet accurate literature discovery fundamentally requires scholarly judgment about what papers actually contribute, a capability that current retrieve-then-rank pipelines cannot provide as they lack mechanisms for deep reasoning over paper content and research context.
Beyond search, existing tools address only isolated aspects of the research process. Citation network visualizations such as Connected Papers [^3] https://www.connectedpapers.com/ reveal paper relationships but lack integrated search and management functions. Reference managers like Zotero [^4] https://www.zotero.org/ and Mendeley [^5] https://www.mendeley.com/ excel at organization yet provide minimal discovery capabilities and require manual curation. Critically, no existing solution bridges the entire research lifecycle from intent-driven discovery through systematic management to continuous frontier tracking, forcing researchers to stitch together fragmented tools with manual effort at each transition point.
We introduce WisPaper, a comprehensive platform that addresses these challenges through three interconnected modules: (1) Scholar Search provides dual-mode retrieval: Quick Search for rapid keyword matching and Deep Search that employs agentic reasoning to evaluate papers’ actual contributions for complex conceptual queries, balancing speed and depth according to user needs. (2) Library enables systematic, custo-
mizable organization of both search results and uploaded documents, supporting long-term knowledge accumulation and continuous updating. (3) AI Feeds automatically recommends highly relevant new papers based on user-defined research interests and library content, enabling researchers to maintain awareness of frontier developments without manual monitoring.
These modules form a closed-loop workflow: researchers search for papers, import and organize them in their Library, while AI Feeds continuously supplements relevant new literature. This integration supports the entire research lifecycle from initial exploration to sustained knowledge management.
Figure 1: WisPaper Knowledge Loop
WisPaper serves diverse user groups including academic researchers, graduate students, and industrial research and development teams across multiple disciplines and languages. By reducing time spent on paper screening and management, the platform enables researchers to allocate more cognitive resources to their core intellectual work. Our contributions are threefold:
(1) A unified platform integrating the entire academic workflow from literature discovery through systematic organization to continuous frontier monitoring, eliminating the need for fragmented tools.
(2) A dual-mode search architecture that combines rapid keyword retrieval with deep semantic reasoning over research context, accommodating diverse search requirements.
(3) A personalized recommendation engine that constructs dynamic user profiles from library curation patterns and explicit interests, reducing manual screening overhead while maintaining research awareness.
2 WisPaper
In this section, we detail the three core modules of WisPaper and their interconnections.
2.1 Scholar Search Module
Scholar Search serves as the primary entry point for literature discovery in WisPaper, offering two complementary search modes that address fundamentally different research scenarios.
2.1.1 Quick Search: Millisecond-Level Responsive Literature Retrieval
Quick Search is is built on efficient keyword retrieval for scenarios where researchers have concrete search targets. When a user has a specific paper title in mind, knows the exact author they want to follow, or needs to quickly scan a well-defined topic, Quick Search provides millisecond-level responses that feel instantaneous.
Each search result presents a complete picture of the paper: beyond the basic title and author information, users can immediately see institutional affiliations, publication dates, venue prestige, citation counts that indicate impact, and abstract snippets that provide context for quick relevance judgments. The interface supports rich interactions—users can access direct citation formatting, express preferences through like and dislike actions, and add papers with available PDFs directly into their Library with one click. More importantly, the system maintains a complete history of all search queries, allowing researchers to re-execute past queries or use them as starting points for refined searches.
2.1.2 Deep Search: Agent-Powered Intelligent Literature Exploration
Deep Search represents a fundamentally different paradigm: an AI agent that conducts intelligent literature exploration on behalf of researchers. Rather than reducing research questions to keywords, the system deploys an autonomous agent (WisModel) that reasons about information needs and orchestrates retrieval tools to discover relevant papers.
Agent-Powered Workflow. The exploration process consists of two key stages orchestrated by WisModel:
  • Query Understanding Stage: The agent interprets the natural language query, performs query decomposition to identify essential aspects, and constructs structured validation criteria (detailed prompt templates are provided in Appendix Table 7).
  • Paper Validation Stage: For each candidate paper, the agent evaluates whether it satisfies the constructed criteria, extracting relevant evidence and producing explicit classifications using a systematic assessment taxonomy (Appendix Table 10). The complete validation prompt structure is shown in Appendix Tables 8 and 9.
Example Workflow. When asked “How do transformer models handle long sequences efficiently?”, WisModel identifies key aspects (transformer architecture, sequence length challenge, efficiency methods), generates targeted search queries, and establishes validation criteria. It then evaluates each retrieved paper against these criteria, classifying results as “Perfect” (meeting all criteria), “Partial” (addressing some aspects), or “No” (failing requirements).
Collaborative Refinement. A distinguishing feature is the agent’s collaborative refinement capability. Researchers can examine and modify the agent-generated criteria to better reflect their information needs, then re-execute the search with refined criteria, progressively narrowing toward their target papers.
2.2 Library Module
Library acts as the central component of WisPaper, evolving the platform from a search tool into a comprehensive knowledge management environment. The design philosophy recognizes that effective research requires structured organization, convenient retrieval, and seamless integration with the reading process.
Organization and Retrieval: Library implements a hierarchical folder system with unlimited nesting depth, allowing researchers to construct taxonomies aligned with their domain structure. Cross-library search addresses the scenario where researchers recall paper content but not storage location, supporting both comprehensive library queries and subtree-constrained searches.
@@ 2.2 Library Module
Paper Import and Metadata Extraction: Papers enter the Library through three mechanisms: direct PDF upload, one-click import from Scholar Search results, or integration from AI Feeds recommendations. Upon import, the system automatically extracts metadata including title, authors, affiliations, abstract, publication venue, date, DOI, and references through PDF structure analysis.
Reading Environment and Ecosystem Integration: The integrated PDF reader extends Library beyond storage into active research support. AI-generated structured summaries enable rapid paper assessment, while community annotations surface peer insights. Personal annotation tools provide highlighting and note-taking with full-text search across notes. Export capabilities support standard bibliographic formats (BibTeX, RIS, EndNote) for manuscript preparation, while synchronization plugins for tools like Zotero and Mendeley enable researchers to maintain WisPaper libraries while using familiar citation workflows.
2.3 AI Feeds Module
AI Feeds addresses one of the most cognitively demanding aspects of active research: staying current with rapidly evolving literature. In fast-moving fields, researchers spend 1-2 hours daily manually scanning journals, conferences, and preprint servers.
@@ 2.3 AI Feeds Module
AI Feeds automates literature monitoring through dual-layer filtering. Users first select broad research areas (e.g., ArXiv categories such as cs.CL for Computation and Language). The system retrieves all papers in these categories since last login, typically 50-100 papers daily per category. Second-layer filtering applies semantic matching against user-defined interests: specific authors, institutions, and topical descriptions. For example, selecting cs.CL with interests in “vision-language models from Google/DeepMind” filters 50-100 daily papers to 3-5 highly relevant matches.
The system learns implicitly from user behavior to refine future filtering. Papers that users add to Library signal strong relevance, while consistently dismissed papers suggest the interest model needs adjustment. This feedback loop operates automatically, gradually improving filtering precision.
2.4 Module Interconnection
WisPaper’s power emerges from the seamless integration of its three modules into a unified research workflow. At the core is a shared paper database that ensures any paper discovered, filtered, or stored in one module is immediately accessible across all others, eliminating redundancy and maintaining consistency.
The modules continuously exchange information to enhance each other’s functionality. Library contents inform AI Feeds’ filtering algorithms, while Scholar Search patterns influence which papers surface in AI Feeds. Papers flow effortlessly between modules with minimal interaction—a single click moves content from discovery to collection while preserving metadata and context.
A typical research session illustrates this integration: check AI Feeds for new publications → add relevant papers to Library → encounter an unfamiliar concept while reading → launch Deep Search from within the reader → organize discovered papers into project folders. Throughout, researchers remain in a single environment without manual data transfer.
3 Deep Search: WisModel Training Pipeline
Technical Implementation Overview. WisPaper’s three modules employ different technical approaches suited to their respective objectives. Quick Search builds upon established keyword-based retrieval systems optimized for millisecond-level response times, while AI Feeds adapts conventional recommendation and filtering frameworks to the academic literature domain. Our primary algorithmic contribution lies in Deep Search’s agent-powered exploration capability, which requires novel training procedures to enable the sophisticated reasoning and validation behaviors described in Section 2.1.2. This section details the training methodology for WisModel, the AI agent that powers Deep Search.
3.1 Training Objectives
WisModel is trained to accomplish two primary objectives that mirror the key stages of agent-powered literature exploration described in Section 2.1.2:
Objective 1: Query Understanding and Criteria Generation.
Given a natural language research query $q$ and current timestamp $t$, the model must generate a set of Boolean search queries $\mathcal{S}=\{s_{1},s_{2},\ldots,s_{k}\}$ (where $k\in[2,4]$) and a set of validation criteria $\mathcal{C}=\{c_{1},c_{2},\ldots,c_{m}\}$ (where $m\in[1,4]$). Each criterion $c_{i}$ consists of a type (e.g., task, methodology, dataset), a descriptive name, a detailed description, and an importance weight $w_{i}$ such that $\Sigma_{i=1}^{m}w_{i}=1$. The complete prompt template is provided in Appendix Table 7.
Objective 2: Paper-Criteria Matching Assessment.
Given a research query $q$, a set of validation criteria $\mathcal{C}$, and paper metadata $p$ (including title, authors, abstract, venue, publication date, and citation count), the model must produce an assessment $a_{i}\in\{\text{support},\text{somewhat\_support},\text{reject},\text{insufficient\_information}\}$ for each criterion $c_{i}\in\mathcal{C}$. The assessment taxonomy is detailed in Appendix Table 10, with prompt templates in Appendix Tables 8 and 9. For each assessment, the model must provide a natural language explanation and extract supporting evidence from the paper metadata.
To provide concrete understanding of the two training objectives, we present two representative examples that demonstrate WisModel’s reasoning process. Table 5 shows a complete query-to-criteria workflow for Objective 1, while Table 6 demonstrates WisModel’s nuanced validation capability for Objective 2—particularly its ability to distinguish between full and partial support, a challenge that proves difficult for baseline models (as quantitatively validated in Section 4.3).
3.2 Training Procedure
Stage 1: Supervised Fine-Tuning.
We initialize WisModel from a pre-trained language model and perform supervised fine-tuning (SFT) on both objectives simultaneously. For Objective 1, the model learns to generate structured JSON outputs containing search queries and criteria. For Objective 2, the model learns to produce detailed assessment objects with criterion-level verdicts, explanations, and evidence extraction. This cold-start phase establishes the basic capability to understand academic queries and reason about paper relevance.
Stage 2: Group Relative Policy Optimization.
Following SFT, we apply Group Relative Policy Optimization (GRPO) [18] specifically to Objective 2 to refine the model’s validation capabilities. Given a prompt $x$ (comprising query, criteria, and paper metadata) with ground truth assessment $a^{*}\in$ {support, somewhat_support, reject, insufficient_information}, the reward function evaluates whether the model’s predicted assessment $a$ aligns with the expert annotation:
$r\text{(}a,a^{*}\text{)}=\mathbb{1}\text{[}a=a^{*}\text{]}=\begin{cases}1&\text{if }a=a^{*}\\ 0&\text{otherwise}\end{cases}$ (1)
GRPO samples $G$ output candidates $\{o_{1},\ldots,o_{G}\}$ from the current policy $\pi_{\theta}$ and computes their respective rewards $\{r_{1},\ldots,r_{G}\}$. We maximize the following objective:
@@ 3.2 Training Procedure
where $A_{i,t}=\text{(}r_{i}-\text{mean}\text{(}\mathbf{r}\text{))}\text{/std}\text{(}\mathbf{r}\text{)}$ is the advantage for each token.

### v3

@@ Abstract
We present WisPaper, an end-to-end agent system that transforms how researchers discover, organize, and track academic literature. The system addresses two fundamental challenges. (1) Semantic search limitations: existing academic search engines match keywords but cannot verify whether papers truly address complex research questions; and (2) Workflow fragmentation: researchers must manually stitch together separate tools for discovery, organization, and monitoring. WisPaper tackles these through three integrated modules. Scholar Search combines rapid keyword retrieval with Deep Search, in which an agentic model, WisModel, validates candidate papers against user queries through structured reasoning. Discovered papers flow seamlessly into Library with one click, where systematic organization progressively builds a user profile that sharpens the recommendations of AI Feeds, which continuously surfaces relevant new publications and in turn guides subsequent exploration, closing the loop from discovery to long-term awareness. On TaxoBench, WisPaper achieves 22.26% recall, surpassing the O3 baseline (20.92%). Furthermore, WisModel attains 93.70% validation accuracy, effectively mitigating retrieval hallucinations.
@@ 1 Introduction
The exponential growth of scientific publications presents unprecedented challenges for researchers attempting to stay current in their fields [article]. Efficiently identifying relevant work and systematically managing findings over time have become increasingly time-consuming and cognitively demanding [kingsley2011]. Two fundamental problems underlie this challenge.
The first is semantic search limitations. Existing academic search engines such as Google Scholar [^1] https://scholar.google.com and Semantic Scholar [^2] https://www.semanticscholar.org/ rely primarily on keyword-based retrieval, which struggles with conceptual queries such as “papers exploring the relationship between in-context learning and inference-time scaling.” A researcher posing this question needs the system to understand what both concepts mean, find papers that engage with their intersection, and judge whether each paper truly contributes—not merely mentions the terms. Recent LLM-based approaches improve query reformulation [jagerman2023queryexpansionpromptinglarge, 10.1145/3539618.3591960, ma-etal-2023-query], yet they still lack mechanisms for deep reasoning over paper content to verify actual relevance [he-etal-2025-pasa].
The second is workflow fragmentation. A typical research workflow has three stages: discovering papers through search engines, organizing them in reference managers like Zotero [^3] https://www.zotero.org/ or Mendeley [^4] https://www.mendeley.com/, and monitoring new publications via arXiv [^5] https://arxiv.org/ or alert services. Each transition requires manual effort: downloading PDFs, copying metadata, tagging entries, checking for updates. Researchers spend significant time on information logistics rather than intellectual work [Bullers2018].
We introduce WisPaper, an end-to-end scholarly agent system that addresses both challenges through three tightly integrated modules (Figure 1). Scholar Search provides dual-mode retrieval: Quick Search for millisecond-level keyword matching, and Deep Search, an agent-powered mode where WisModel, a specialized LLM trained via reinforcement learning, decomposes queries into verification criteria and validates papers through structured reasoning. Discovered papers flow seamlessly into Library with one click, eliminating manual downloading and metadata entry. As papers accumulate, Library progressively builds a user profile that sharpens the recommendations of AI Feeds, which continuously surfaces relevant new publications and in turn guides subsequent exploration. Over time, each iteration of the cycle makes the next more efficient: collections grow richer, profiles become more precise, and recommendations align more closely with evolving research interests.
Figure 1: The WisPaper knowledge loop. Researchers discover papers via Scholar Search, organize them in Library, and receive continuous updates through AI Feeds. Modules share a unified paper database and exchange information to enhance each other.
Our main contributions are as follows:
(1) We present WisPaper, an end-to-end scholarly agent system in which search, organization, and monitoring form a closed knowledge loop.
(2) We develop WisModel, an agentic model trained via reinforcement learning that decomposes complex research queries into editable verification criteria and validates candidate papers through structured reasoning.
(3) We demonstrate that WisPaper achieves 22.26% recall on the TaxoBench benchmark and 93.70% accuracy in criteria matching, surpassing existing state-of-the-art agents.
Figure 2: Deep Search workflow. In Stage 1, WisModel decomposes the user’s natural language query into structured, executable queries and validation criteria. In Stage 2, the model validates candidate papers, strictly requiring direct quotes from the original text as evidence to eliminate hallucinations. Users can collaboratively refine criteria and re-execute the search.
2 The WisPaper Knowledge Loop
WisPaper unifies literature discovery, organization, and monitoring into a single platform whose modules continuously exchange information to enhance each other, forming a knowledge loop.
A Typical Session.
A researcher’s daily workflow illustrates the knowledge loop: (1) Open AI Feeds to review 3–5 papers recommended overnight (§2.3). (2) Add interesting papers to Library with one click; the system auto-extracts metadata (§2.2). (3) While reading, encounter an unfamiliar technique; launch Deep Search directly from the PDF reader (§2.1). (4) Import newly found papers into Library, further refining the profile and improving tomorrow’s AI Feeds. Sections 2.12.3 detail each module.
2.1 Scholar Search
Scholar Search offers two complementary modes for literature discovery.
Quick Search.
For well-defined information needs (e.g., looking up a specific paper by title or author), Quick Search provides millisecond-level keyword retrieval. Each result card displays title, authors, affiliations, venue, citation count, and an abstract snippet for rapid relevance judgments. Users can seamlessly import papers to Library with one click, copy formatted citations, and express relevance preferences (like/dislike) that feed into the user profile.
Deep Search.
Deep Search targets intent-heavy queries that cannot be reliably answered by keyword matching (e.g., “find recent AI4Science papers that evaluate model generalization on real lab data”). It invokes WisModel (Section 3) to run a two-stage pipeline that is transparent and editable:
Stage 1: Intent Decomposition. WisModel first expands the user query into multiple distinct search Queries to quickly retrieve a high-recall set of potentially relevant paper candidates. It then constructs an explicit verification checklist that specifies what must hold for a paper to qualify (e.g., domain scope, method constraints). Both the search queries and the verification checklist are presented to the user as editable UI controls.
Stage 2: Evidence-Backed Verification. For each retrieved candidate, WisModel produces a structured report that links each checklist item to concrete evidence from the paper (e.g., title, abstract, and metadata) and an assessment label. Results are ranked by the aggregated checklist satisfaction, and the full per-criterion trace is exposed to support rapid inspection, thereby making each inclusion decision transparent and evidence-grounded.
Collaborative Refinement. Because the search queries and checklist are editable, users can revise retrieval terms and verification criteria and re-run verification without rephrasing the original query, progressively steering the search while keeping the system’s reasoning auditable.
Seeding the Knowledge Loop.
Discovered papers flow directly into Library: one-click paper import transfers full metadata to construct a personalized knowledge base, while simultaneously seeding the interest signals that refine the user profile within Library (§2.2).
2.2 Library
Library turns WisPaper from a search tool into a personal knowledge hub for papers. It streamlines paper ingestion and organization while continuously updating a user interest profile that drives the knowledge loop with AI Feeds.
Paper Import and Metadata Extraction.
@@ 2.2 Library
Papers enter Library through three pathways: direct PDF upload, one-click import from Scholar Search and AI Feeds. Upon ingestion, the system automatically extracts structured metadata (title, authors, affiliations, abstract, venue, date, DOI, and references) via PDF structure analysis, eliminating the manual entry that plagues traditional reference managers.
Organization and Export.
Library implements a hierarchical folder system with unlimited nesting (e.g., “Survey $\rightarrow$ Retrieval $\rightarrow$ Dense Retrieval”), and supports both full-collection and subtree-constrained search, allowing users to locate papers by content rather than just by storage path. For seamless interoperability, it exports to standard bibliographic formats (BibTeX, RIS, EndNote) and offers synchronization plugins for Zotero and Mendeley.
User Interest Profiling.
Library maintains an evolving interest profile that summarizes what the user is currently reading and collecting, which is subsequently consumed by AI Feeds (§2.3). The profile is updated from both explicit signals (e.g., user-specified interests and curated folders) and implicit behavioral signals (e.g., imported papers and how they are organized). Concretely, each newly added paper contributes topical/entity cues derived from its metadata (title, abstract, keywords, venue, authors), while the assigned folder path provides fine-grained, user-defined labels that disambiguate subtopics.
Driving the Knowledge Loop.
As papers accumulate and are organized, these import patterns and assigned folder hierarchies build a dynamic user interest profile that drives AI Feeds, making tomorrow’s recommendations more precisely targeted.
2.3 AI Feeds
In fast-moving fields, researchers may spend 1–2 hours daily scanning journals and preprint servers. AI Feeds turns this passive monitoring into an active, personalized process through a dual-layer pipeline that dynamically adapts to the evolving user interests.
Dual-Layer Filtering.
@@ 2.3 AI Feeds
The first layer handles source selection: users subscribe to broad research areas (e.g., ArXiv categories such as cs.CL), and the system retrieves all new papers published since the last login, typically 50–100 per category per day. The second layer applies personalized filtering by matching these candidates against the user’s current interest profile, narrowing daily papers to 3–5 highly relevant matches. Initially seeded by user-defined interests (e.g., “vision-language models from Google/DeepMind”), this baseline profile is subsequently and continuously enriched by the Library module (§2.2) as the user interacts with the system.
Completing the Knowledge Loop.
AI Feeds completes the loop by channeling attention back into discovery: a recommended paper can be saved to Library with one click or can spark a new Scholar Search (§2.1), restarting the cycle with sharpened focus.
3 WisModel: The Deep Search Agent
Unlike standard conversational assistants that tolerate loose information retrieval, a scholarly agent faces a critical evidence-grounding challenge: it must translate underspecified user intents into precise epistemological criteria and execute rigorous, hallucination-free validation over complex academic texts. Standard LLMs often fail at this, prone to ignoring nuanced constraints or fabricating supporting evidence. To overcome this, we formulate Deep Search as a structured reasoning task and train WisModel via a bespoke Reinforcement Learning (RL) pipeline to enforce both logical rigor and structural adherence.
3.1 Task Formulation & Initialization
We formalize the agent’s reasoning process into two sequential objectives:
Objective 1: Intent Decomposition.
Given a raw user query $q$, the model generates a retrieval plan consisting of search queries $\mathcal{S}$ and a set of weighted verification criteria $\mathcal{C}=\{(c_{i},w_{i})\}_{i=1}^{m}$, where each $c_{i}$ represents a criteria and $w_{i}$ denotes its weight. Table 5 provides a complete demonstration example of intent decomposition. The complete prompt template is provided in Table 7.
Objective 2: Evidence-Backed Validation.
Given $q$, $\mathcal{C}$, and paper metadata $p$, the model acts as a rigorous reviewer, producing a structured verification report $y$. For each criterion $c_{i}$, the model must output a tuple $(a_{i},e_{i},r_{i})$:
  • Assessment $a_{i}\in$ {support, somewhat support, reject, insufficient information}: A discrete logical verdict. The assessment taxonomy is shown in Table 10, with prompt templates in Tables 8 and 9.
  • Evidence Snippet $e_{i}$: A verbatim quote extracted from $p$ that justifies $a_{i}$, strictly prohibiting hallucinated text.
  • Rationale $r_{i}$: A concise natural language explanation connecting $e_{i}$ to $c_{i}$.
We initialize WisModel via Supervised Fine-Tuning (SFT) on an expert-annotated dataset. This stage bootstraps the model’s ability to follow the complex JSON schema and understand the fundamental mapping between academic discourse and verification criteria.
3.2 Reinforcement Learning via Multi-Dimensional Shaped Rewards
While SFT establishes baseline capabilities, it often struggles to penalize subtle logical contradictions or hallucinated evidence—a common failure mode in academic verification. To address this, we further optimize the policy $\pi_{\theta}$ using Group Relative Policy Optimization (GRPO). GRPO samples $G$ output candidates $\{o_{1},\ldots,o_{G}\}$ from the current policy $\pi_{\theta}$ and computes their respective rewards $\{r_{1},\ldots,r_{G}\}$. We maximize the following objective:
@@ 3.2 Reinforcement Learning via Multi-Dimensional Shaped Rewards
where $A_{i,t}=\text{(}r_{i}-\text{mean}\text{(}\mathbf{r}\text{))}\text{/std}\text{(}\mathbf{r}\text{)}$ is the advantage for each token. Instead of a sparse binary reward, we design a Multi-Dimensional Shaped Reward function $\mathcal{R}(y)$ that decomposes the verification quality into four orthogonal dimensions:
  1. 1.
    Format Consistency ($r_{\text{fmt}}$): Ensures deterministic system execution by rewarding strict adherence to the nested JSON schema and penalizing parsing errors.
  2. 2.
    Faithful Grounding ($r_{\text{ground}}$): Specifically targets hallucination. It computes the longest common subsequence between the generated evidence $e_{i}$ and the source text $p$. If $e_{i}$ is not a substring of $p$, a severe penalty is applied.
  3. 3.
    Logical Entailment ($r_{\text{logic}}$): Rewards the correctness of the discrete verdict $a_{i}$ against expert labels, heavily penalizing contradictions (e.g., predicting “support” when the paper explicitly “rejects”).
  4. 4.
    Reasoning Alignment ($r_{\text{reason}}$): Measures the semantic similarity of the generated rationale $r_{i}$ to expert explanations using embedding distance, ensuring the reasoning path is sound.
Dynamic Curriculum Optimization.
Directly optimizing all rewards can lead to instability. We explicitly adopt a curriculum strategy by introducing dynamic time-dependent weights $\lambda_{t}$ for each reward component. The total reward at training step $t$ is defined as:
$\mathcal{R}_{t}\text{(}y\text{)}=\lambda_{t}^{\text{fmt}}r_{\text{fmt}}\text{+}\lambda_{t}^{\text{ground}}r_{\text{ground}}\text{+}\lambda_{t}^{\text{sem}}\text{(}r_{\text{logic}}\text{+}r_{\text{reason}}\text{)}$ (2)
Early in training, we prioritize $\lambda^{\text{fmt}}$ to enforce structural validity. As the KL divergence stabilizes, we decay $\lambda^{\text{fmt}}$ and increase $\lambda^{\text{ground}}$ and $\lambda^{\text{sem}}$, shifting the optimization focus from syntax to deep semantic reasoning. GRPO then updates the policy by favoring higher-reward trajectories relative to the group mean, effectively aligning the model with the rigorous standards of scientific inquiry.